X/2-7=-6 (Please check your answer)
Answer by checkley79(3341)

X/2-7=-6
x/2=-6+7
x/2=1 cross multiply.
x=1*2=x=2 ans.
Proof:
2/2-7=-6
1-7=-6
-6=-6